OpenStudy (jazzyluvsu123):

How do i convert 5,184 square inches to Square feet would i Multiply or Divide? Medal

OpenStudy (anonymous):

divide it by 12

OpenStudy (jazzyluvsu123):

Then to Convert Square feet to Square yards i divide by 3 ?

OpenStudy (welshfella):

divide by 12*12

OpenStudy (anonymous):

yes

OpenStudy (jazzyluvsu123):

So Converting 5,184 Square inches to Square yards my answer would be - 144

OpenStudy (anonymous):

yes, see that wasnt so hard

OpenStudy (jazzyluvsu123):

OMG. I typed in the answer and it said i was wrong. the answer was 4 ????

OpenStudy (jazzyluvsu123):

im so confused?

OpenStudy (jazzyluvsu123):

Converting 5,184 Square Inches into Square yards? the answer was 4 ?

OpenStudy (welshfella):

1 square yard = 36*36 = 1296 sq ins 5184 sq ins = 5184 / 1296 = 4 square yards

OpenStudy (jazzyluvsu123):

Thanks @Welshfella